Excess hydrochloric acid is reacted with a sample of potassium carbonate. After drying, the potassium chloride weighs 2.31 g. What is the volume of carbon dioxide gas produced if the reaction occurs at 21 C and 746 mmHg?

K2CO3 (s) + 2HCI (aq) --------> 2KCI (aq) + H2O (I) + CO2 (g)

A. 0.381 L

B. 0.192 L

C. 1.52 L

D. 0.757 L
Answer #1

molar mass of KCl = 74.6 g/mol
number of mole of KCl = (mass of KCl)/(molar mass of KCl)
= 2.31/74.6
= 0.0310 mole

according to cofficient of balanced equation
number of mole of CO2 = (number of mole of KCl)/2
= 0.0310/2
= 0.0155 mole

for CO2 use,
P*V = (number of mole of CO2)*R*T
T = 21 oC
= 294 k
P = 746 mmHg
= 0.982 atm
so,
0.982*V = 0.0155*0.082*294
V = 0.381 L

Answer : 0.381 L